OpenStudy (anonymous):

3 letters are dictated to 3 persons and an envelope is addressed to each of them,the letters are inserted into the envelopes at random so that each envelope contains exactly 1 letter.Find the probability that at least 1 letter is in its proper envelope?

OpenStudy (anonymous):

1-1/(3*2*1)=5/6

OpenStudy (anonymous):

the answer given is 2/3..

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

It is 1- 2/(3*2*1) =1-1/3=2/3

OpenStudy (anonymous):

probability that at least 1 letter is in its proper envelope= 1- probability that none of the letter is in its proper envelop= 1- 2/(3*2*1) =2/3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@sauravshakya how did u get the probabily of geeting none?

OpenStudy (anonymous):

No. of ways that none of the letter goes in its proper=2 Total no. of ways letter can be put in the enelope=3*2*1

OpenStudy (anonymous):

@sauravshakya thank u...

OpenStudy (anonymous):

welcome